A triangle plot of land has a perimeter of 1170 metre.
The first side is 240 metres less than twice the second side.
the third side is half the combined length of the first and second side

Answer by mananth(16949)

A triangle plot of land has a perimeter of 1170 metre.
The first side is 240 metres less than twice the second side.
the third side is half the combined length of the first and second side
...
second side=x
first side = 2x-240
third side=1/2*(x+2x-240)
...
x+2x-240+1/2*(3x-240)=1170
...
multiply the equation by 2
2x+4x-480+3x-240=1170*2
...
9x-720=2340
+720 +720
9x=2340+720
9x=3060
/9
x=340 meters the second side
..
first side = 2x-240=680-240=440 meters
...
third side = 390 meters
...
m.ananth@hotmail.ca